### v1

# A classification of incompleteness statements

Abstract.
For which choices of $X,Y,Z\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$ does no sufficiently strong $X$-sound and $Y$-definable extension theory prove its own $Z$-soundness? We give a complete answer, thereby delimiting the generalizations of Gödel’s second incompleteness theorem that hold within second-order arithmetic.
Thanks to Hanul Jeon for discussion. The first author was partially supported by NSF Grant DMS-2054379.
1. Introduction
Gödel’s second incompleteness theorem states that no sufficiently strong consistent and recursively axiomatized theory proves its own consistency. We give an equivalent restatement here:
Theorem 1.1 (Gödel).
No sufficiently strong $\Pi^{0}_{1}$-sound and $\Sigma^{0}_{1}$-definable theory proves its own $\Pi^{0}_{1}$-soundness.
A recent result [3] lifts Gödel’s theorem to the setting of second-order arithmetic, where stronger reflection principles are formalizable:
Theorem 1.2 (Walsh).
No sufficiently strong $\Pi^{1}_{1}$-sound and $\Sigma^{1}_{1}$-definable theory proves its own $\Pi^{1}_{1}$-soundness.
This latter theorem applies to $\Sigma^{1}_{1}$-definable theories—not just to the narrower class of $\Sigma^{0}_{1}$-definable theories—whence it is not an immediate consequence of Gödel’s original theorem.
There are three classes of formulas in the statement of Theorem 1.2, leading to eight variations one could consider, including the original. In this paper we consider the other seven. The following table records the truth-values of the statement: No sufficiently strong $X$-sound and $Y$-definable theory proves its own $Z$-soundness.
X Y Z
$\Pi^{1}_{1}$ $\Sigma^{1}_{1}$ $\Pi^{1}_{1}$
$\Pi^{1}_{1}$ $\Pi^{1}_{1}$ $\Pi^{1}_{1}$
$\Sigma^{1}_{1}$ $\Pi^{1}_{1}$ $\Pi^{1}_{1}$
$\Pi^{1}_{1}$ $\Sigma^{1}_{1}$ $\Sigma^{1}_{1}$
$\Sigma^{1}_{1}$ $\Sigma^{1}_{1}$ $\Sigma^{1}_{1}$
$\Pi^{1}_{1}$ $\Pi^{1}_{1}$ $\Sigma^{1}_{1}$
$\Sigma^{1}_{1}$ $\Sigma^{1}_{1}$ $\Pi^{1}_{1}$
$\Sigma^{1}_{1}$ $\Pi^{1}_{1}$ $\Sigma^{1}_{1}$
To place the ✗s on the table we show how to give appropriately non-standard definitions of arbitrarily strong sound theories. Theorem 1.2 places the first ✓ on the table; for this a “sufficiently strong” theory is any extension of $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$. To place the second ✓ on the table we forge a connection between provable $\Sigma^{1}_{1}$-soundness and a kind of “pseduo-ordinal analysis.” Whereas $\Pi^{1}_{1}$-soundness provably follows from the well-foundedness of a theory’s proof-theoretic ordinal, we show that $\Sigma^{1}_{1}$-soundness provably follows from the statement that a certain canonical ill-founded linear order lacks hyperarithmetic descending sequences. For the second ✓ a “sufficiently strong” theory is any extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$.
2. The Proofs
2.1. Simplest Cases
We begin by placing the first four ✗s on the table.
When $\Gamma$ is a set of formulas, we write $\mathsf{RFN}_{\Gamma}(U)$ for the sentence stating the $\Gamma$-soundness of $U$ (i.e. reflection for formulas from $\Gamma$):
$\mathsf{RFN}_{\Gamma}(U):=\forall\varphi\in\Gamma\big{(}\mathsf{Pr}_{U}(\varphi)\to\mathsf{True}_{\Gamma}(\varphi)\big{)}.$
Theorem 2.1.
Let $\Gamma\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$. For any sound and arithmetically definable theory $S$, there is a sound and $\Gamma$-definable extension $T$ of $S$ such that $T\vdash\mathsf{RFN}_{\Gamma}(T)$.
Proof.
We define $U:=S+\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$. Then we define:
$T(\varphi):=U(\varphi)\wedge\mathsf{RFN}_{\Gamma}(U)$
That is, $\varphi\in T$ if and only if both $\varphi\in U$ and $\mathsf{RFN}_{\Gamma}(U)$.
Then $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}\vdash T=\emptyset\vee\big{(}T=U\wedge\mathsf{RFN}_{\Gamma}(U)\big{)}$. Thus, reasoning by cases, $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}\vdash\mathsf{RFN}_{\Gamma}(T)$. Since $T=U\supseteq\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$, $T\vdash\mathsf{RFN}_{\Gamma}(T)$.
To see that $T$ is $\Gamma$-definable, note that $U$ is $\Gamma$-definable and that $\mathsf{RFN}_{\Gamma}(U)$ has an arithmetic antecedent and a $\Gamma$ consequent.
Finally, note that $T$ is just $U$, whence it is sound. ∎
Remark 2.2.
In the proof of Theorem 2.1, we use the $\Sigma^{1}_{1}$ choice principle only if $\Gamma=\Sigma^{1}_{1}$. Indeed, to infer that $\mathsf{RFN}_{\Sigma^{1}_{1}}(U)$ is $\Sigma^{1}_{1}$, we must pull the positively occurring existential set quantifier from $\mathsf{True}_{\Gamma}(\varphi)$ in front of a universal number quantifier. If $\Gamma=\Pi^{1}_{1}$, it suffices to define $U$ as $S+\mathsf{ACA}_{0}$, since $\mathsf{RFN}_{\Pi^{1}_{1}}$ has a finite axiomatization in $\mathsf{ACA}_{0}$.
2.2. Intermediate Cases
We can resolve two more cases with a subtler version of the proof of Theorem 2.1. First, we recall the following useful lemma. For $\Gamma\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$, we let $\widehat{\Gamma}$ be the dual complexity class.
Lemma 2.3.
For $T$ extending $\mathsf{ACA}_{0}$, $\mathsf{RFN}_{\widehat{\Gamma}}(T)$ does not follow from any consistent extension of $T$ by $\Gamma$ formulas.
Proof.
Suppose $T+\gamma\vdash\mathsf{RFN}_{\widehat{\Gamma}}(T)$ with $\gamma\in\Gamma$. Then $T+\gamma\vdash\mathsf{Pr}_{T}(\neg\gamma)\to\neg\gamma$. Hence $T+\gamma\vdash\neg\mathsf{Pr}_{T}(\neg\gamma)$, i.e., $T+\gamma\vdash\mathsf{Con}(T+\gamma)$. So $T+\gamma\vdash\bot$. ∎
The following theorem adds two more ✗s to our table.
Theorem 2.4.
Let $\Gamma\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$. For any sound and arithmetically definable theory $U$, there is a $\widehat{\Gamma}$-sound and $\widehat{\Gamma}$-definable extension of $U$ that proves its own $\Gamma$-soundness.
Proof.
Consider the following formulas:
$\varphi(x)$ $:=x=\ulcorner\mathsf{RFN}_{\Gamma}(U)\urcorner\vee x=\ulcorner\neg\mathsf{RFN}_{\widehat{\Gamma}}(U+\mathsf{RFN}_{\Gamma}(U))\urcorner$
$\tau(x)$ $:=x=U(x)\vee\Big{(}\mathsf{RFN}_{\widehat{\Gamma}}\big{(}U+\mathsf{RFN}_{\Gamma}(U)\big{)}\wedge\varphi(x)\Big{)}$
Let $T$ be the theory defined by $\tau$.
Claim.
$T$ is $\widehat{\Gamma}$-definable via $\tau$.
By inspection.
Claim.
$T$ is $\widehat{\Gamma}$-sound.
Since $U$ is sound, $U+\mathsf{RFN}_{\Gamma}(U)$ is sound, so $\mathsf{RFN}_{\widehat{\Gamma}}(U+\mathsf{RFN}_{\Gamma}(U))$ holds, and therefore externally, we see that $T$ is the theory:
$U+\mathsf{RFN}_{\Gamma}(U)+\neg\mathsf{RFN}_{\widehat{\Gamma}}(U+\mathsf{RFN}_{\Gamma}(U)).$
In particular, $T$ has the form $U^{\prime}+\neg\mathsf{RFN}_{\widehat{\Gamma}}(U^{\prime})$ where $U^{\prime}$ is sound. Suppose that $U^{\prime}+\neg\mathsf{RFN}_{\widehat{\Gamma}}(U^{\prime})\vdash\sigma$ where $\sigma$ is false $\widehat{\Gamma}$. Then $U^{\prime}+\neg\sigma\vdash\mathsf{RFN}_{\widehat{\Gamma}}(U^{\prime})$. So $\mathsf{RFN}_{\widehat{\Gamma}}(U^{\prime})$ follows from a consistent extension of $U^{\prime}$ by $\Gamma$ formulas, contradicting Lemma 2.3.
Claim.
$T\vdash\mathsf{RFN}_{\Gamma}(\tau)$.
From our external characterization of $T$ we see that
$T\vdash\neg\mathsf{RFN}_{\widehat{\Gamma}}(U+\mathsf{RFN}_{\Gamma}(U)).$
Hence $T$ proves that $\tau$ defines the theory $U$. Again, appealing to our external characterization of $T$, $T\vdash\mathsf{RFN}_{\Gamma}(U)$. Thus, $T\vdash\mathsf{RFN}_{\Gamma}(\tau)$. ∎
2.3. Limitations
The presentation $\tau$ of theory $T$ defined in Theorem 2.4 is clearly somewhat pathological, in part because $T$ cannot discern the identity of $\tau$. Before continuing to the final case, we want to illustrate that such pathologies are inevitable. We use a proof technique suggested at the end of [3].
Proposition 2.5.
Let $T$ be a $\Gamma$-definable extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ that proves Theorem 1.2 and Theorem 2.7. Suppose that there is a $\Gamma$ presentation $\tau$ of $T$ such that $T$ proves $\mathsf{RFN}_{\widehat{\Gamma}}(\tau)$. Then both of the following hold:
  1. (1)
    There is a theorem $A$ of $T$ such that $T\vdash\neg\tau(A)$.
  2. (2)
    There is a $\Gamma$ presentation $\tau^{\star}$ of $T$ such that $T$ proves $\neg\mathsf{RFN}_{\widehat{\Gamma}}(\tau^{\star})$.
Proof.
Suppose that each of the following holds:
  1. (1)
    $T$ is definable by a $\Gamma$ formula $\tau$;
  2. (2)
    $T$ extends $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$;
  3. (3)
    $T$ proves Theorem 1.2 and Theorem 2.7;
  4. (4)
    $T$ proves the $\widehat{\Gamma}$-soundness of $\tau$.
Let $\sigma$ be a sentence axiomatizing $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$. We have assumed $T\vdash\sigma$. We also have that $T\vdash\mathsf{RFN}_{\widehat{\Gamma}}(\tau)$. Let $A_{1},\dots,A_{n}$ be the axioms of $T$ that are used in the $T$-proof of $\sigma\wedge\mathsf{RFN}_{\widehat{\Gamma}}(\tau)$. Thus:
$\vdash(A_{1}\wedge\dots\wedge A_{n})\to\big{(}\sigma\wedge\mathsf{RFN}_{\widehat{\Gamma}}(\tau)\big{)}.$
Claim.
$T\vdash\tau(A_{1}\wedge\dots\wedge A_{n})\to\neg\mathsf{RFN}_{\widehat{\Gamma}}(\tau).$
Reason in $T$. Suppose $\tau(A_{1}\wedge\dots\wedge A_{n})$. Then $\tau$ extends $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ and $\tau$ proves $\mathsf{RFN}_{\widehat{\Gamma}}(\tau)$. Since $\tau$ is a $\Gamma$ formula, Theorem 1.2 (if $\Gamma=\Sigma^{1}_{1}$) or Theorem 2.7 (if $\Gamma=\Pi^{1}_{1}$) entails that $\tau$ is not $\widehat{\Gamma}$-sound.
Since $T\vdash\mathsf{RFN}_{\widehat{\Gamma}}(\tau)$, the claim implies that $T\vdash\neg\tau(A_{1}\wedge\dots\wedge A_{n})$.
On the other hand, consider $\tau^{\star}(x):=\tau(x)\vee x=\ulcorner A_{1}\wedge\dots\wedge A_{n}\urcorner$. Note that $\tau^{\star}$ is a $\Gamma$ definition of $T$. Yet we have just shown that $T\vdash\neg\mathsf{RFN}_{\widehat{\Gamma}}(\tau^{\star})$. ∎
Remark 2.6.
Note that in the proof we need only assume that $T$ extends $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ if $\Gamma=\Pi^{1}_{1}$. If $\Gamma=\Sigma^{1}_{1}$, it suffices to assume that $T$ extends $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$ since Theorem 1.2 applies to extensions of $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$. Likewise, we need not assume that $T$ proves both Theorem 1.2 and Theorem 2.7. It suffices to assume that $T$ proves Theorem 1.2 (if $\Gamma=\Sigma^{1}_{1}$) or that $T$ proves Theorem 2.7 (if $\Gamma=\Pi^{1}_{1}$).
2.4. Hardest Case
The only remaining case is the dual form of Theorem 1.2:
Theorem 2.7.
No $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ proves its own $\Sigma^{1}_{1}$-soundness.
Theorem 1.2 was proved using concepts from ordinal analysis. In short, a connection is forged between $\Pi^{1}_{1}$-soundness and well-foundedness of proof-theoretic ordinals. Since we are now interested in $\Sigma^{1}_{1}$-soundness, we forge an analogous connection between $\Sigma^{1}_{1}$-soundness and pseudo-well-foundedness, where an order is pseudo-well-founded if it lacks hyperarithmetic descending sequences.
For the rest of this section assume that $T$ is a $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$. In what follows, $\mathsf{PWF}(x)$ is a predicate stating that $x$ encodes a recursive pseudo-well-founded order (that is, a linear order with no hyperarithmetic decreasing sequence). Note that in $\mathsf{ATR}_{0}$, a subsystem of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$, a universal quantifier over Hyp can be transformed into an existential set quantifier. This is [2, Theorem VIII.3.27] (originally proved by Spector and Gandy):
Theorem 2.8.
Let $\varphi(X)$ be a $\Sigma^{1}_{1}$ formula with no free set variables other than $X$. Then we can find an arithmetical formula $\theta(X,Z)$ such that
$\mathsf{ATR}_{0}\vdash\forall X\big{(}\varphi(X)\leftrightarrow\forall Z\leq_{H}X\;\theta(X,Z)\big{)}.$
Since $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ contains $\mathsf{ATR}_{0}$, the statement $\mathsf{PWF}(x)$ is $T$-provably equivalent to a $\Sigma^{1}_{1}$ formula.
We will define $\prec_{T}$ to hold on pairs $(e,\alpha)$ where $e\in\textsc{Rec}$ and $\alpha\in dom(\prec_{e})$. We define $(e,\alpha)\prec_{T}(e^{\prime},\beta)$ to hold
if there is some $f\in\textsc{Hyp}$ so that $\mathsf{Emb}(f,\prec_{e}\upharpoonright\alpha+1,\prec_{e^{\prime}}\upharpoonright\beta)$ and $\mathsf{Pr}_{T}(\mathsf{PWF}(\prec_{e^{\prime}}))$.
To prove that $T\nvdash\mathsf{RFN}_{\Sigma^{1}_{1}}(T)$ it suffices to check that $T\vdash\mathsf{RFN}_{\Sigma^{1}_{1}}(T)\to\mathsf{PWF}(\prec_{T})$ and that $T\nvdash\mathsf{PWF}(\prec_{T})$. Let’s take these one at a time.
Claim.
$T\vdash\mathsf{RFN}_{\Sigma^{1}_{1}}(T)\to\mathsf{PWF}(\prec_{T})$.
Proof.
Reason in $T$. Suppose $\neg\mathsf{PWF}(\prec_{T})$. That is, there is a hyp descending sequence $f$ in $\prec_{T}$. Let $f(n)=(e_{n},\beta_{n})$. Thus we have:
$\forall n\;(e_{n+1},\beta_{n+1})\prec_{T}(e_{n},\beta_{n})$
By the definition of $\prec_{T}$, this is just to say:
$\forall n\;\exists f\in\textsc{Hyp}\;\mathsf{Emb}\Big{(}(f,\prec_{e_{n}}\upharpoonright\alpha+1,\prec_{e_{n+1}}\upharpoonright\beta)\wedge\mathsf{Pr}_{T}\big{(}\mathsf{PWF}(\prec_{e_{n+1}})\big{)}\Big{)}.$
Let’s rewrite this:
$\forall n\;\exists f\Bigg{(}f\in\textsc{Hyp}\wedge\mathsf{Emb}\Big{(}(f,\prec_{e_{n}}\upharpoonright\alpha+1,\prec_{e_{n+1}}\upharpoonright\beta)\wedge\mathsf{Pr}_{T}\big{(}\mathsf{PWF}(\prec_{e_{n+1}})\big{)}\Big{)}\Bigg{.}$
Note that the predicate $f\in\textsc{Hyp}$ is $\Delta^{1}_{1}$ in $\mathsf{ACA}_{0}$ via the Kleene–Souslin Theorem [2, Theorem VIII.3.19]. Since $T$ is $\Pi^{1}_{1}$-definable, this means that the matrix of the formula above is generalized $\Pi^{1}_{1}$. Using choice for generalized $\Pi^{1}_{1}$ formulas (this is an application of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$), we may swap the quantifiers in the above:
$\exists f\;\forall n\Big{(}f_{n}\in\textsc{Hyp}\wedge\mathsf{Emb}(f_{n},\prec_{e_{n}}\upharpoonright\alpha+1,\prec_{e_{n+1}}\upharpoonright\beta)\wedge\mathsf{Pr}_{T}\big{(}\mathsf{PWF}(\prec_{e_{n+1}})\big{)}\Big{)}.$
Note that the witness $f$ to this formula is technically a set that encodes countably many functions $(f_{n})_{n<\omega}$ in the standard way.
Using arithmetic comprehension, we form the composition $f_{\star}$ of the functions encoded in $f$. The function $f_{\star}$ is a hyp descending sequence in $\prec_{e_{0}}$. So $\prec_{e_{0}}$ is pseudo-ill-founded even though $T$ proves the opposite. Hence, $\neg\mathsf{RFN}_{\Sigma^{1}_{1}}(T)$. ∎
Before addressing the second claim, let’s record a dual form of Rathjen’s formalized version of $\Sigma^{1}_{1}$ bounding [1, Lemma 1.1].
Lemma 2.9.
Suppose $H(x)$ is a $\Pi^{1}_{1}$ formula such that
$\mathsf{ATR}_{0}\vdash\forall x\big{(}H(x)\to\mathsf{PWF}(x)\big{)}.$
Then for some $e\in\textsc{Rec}$, $\mathsf{ATR}_{0}\vdash\mathsf{PWF}(e)\wedge\neg H(e)$.
Proof.
Since $\mathsf{PWF}$ (the set of pseduo-well-founded recursive linear orders) is $\Sigma^{1}_{1}$-complete, there is a total recursive function $\{k\}$ such that:
$\neg H(n)\Longleftrightarrow\mathsf{PWF}\big{(}\{{k}\}(n)\big{)}.$
The reduction can be carried out in $\mathsf{ATR}_{0}$, so
$\mathsf{ATR}_{0}\vdash\neg H(x)\leftrightarrow\mathsf{PWF}\big{(}\{{k}\}(x)\big{)}.$
By the recurision theorem and the S-m-n theorem, there is an integer $e$ so that $\mathsf{ATR}_{0}$ proves that $\forall i[\{e\}(i)\simeq\{\{k\}(e)\}(i)]$ (where $\simeq$ means that if either side converges then both sides converge and are equal). Working in $\mathsf{ATR}_{0}$, $\neg\mathsf{PWF}(e)$ implies $\neg\mathsf{PWF}(\{k\}(e))$, which implies $H(e)$, which implies $\mathsf{PWF}(e)$, which is a contradiction. So $\mathsf{ATR}_{0}\vdash\mathsf{PWF}(e)$. (Not that this implies $e\in\textsc{Rec}$ by the definition of $\mathsf{PWF}(e)$.)
Similarly, $H(e)$ implies $\neg\mathsf{PWF}(\{k\}(e))$, which is equivalent to $\neg\mathsf{PWF}(e)$, which we have already ruled out. So $\mathsf{ATR}_{0}\vdash\neg H(e)$. ∎
Claim.
$T\nvdash\mathsf{PWF}(\prec_{T})$.
Proof.
Suppose that $T$ proves $\mathsf{PWF}(\prec_{T})$. From the definition of $\prec_{T}$, it follows that:
$T\vdash\big{(}\exists f\in\textsc{Hyp}\;\mathsf{Emb}(f,\prec_{x},\prec_{T})\big{)}\to\mathsf{PWF}(\prec_{x}).$
By Theorem 2.8 there exists a $\Pi^{1}_{1}$ formula $\pi(x)$ such that:
$\mathsf{ATR}_{0}\vdash\pi(x)\leftrightarrow\exists f\in\textsc{Hyp}\;\mathsf{Emb}(f,\prec_{x},\prec_{T}).$
By Lemma 2.9, there is some $e$ so that
$\mathsf{ATR}_{0}\vdash\mathsf{PWF}(\prec_{e})\wedge\neg\pi(e).$
Hence $\mathsf{ATR}_{0}\vdash\neg\exists f\in\textsc{Hyp}\;\mathsf{Emb}(f,\prec_{e},\prec_{T})$. Moreover, since $\mathsf{ATR}_{0}$ is sound, we infer that $\neg\exists f\in\textsc{Hyp}\;\mathsf{Emb}(f,\prec_{e},\prec_{T})$ is true.
On the other hand, since $T$ extends $\mathsf{ATR}_{0}$, we infer that $T\vdash\mathsf{PWF}(\prec_{e})$. Hence the map $\alpha\mapsto(e,\alpha)$ is a canonical hyp embedding of $\prec_{e}$ into $\prec_{T}$. So $\neg\exists f\in\textsc{Hyp}\;\mathsf{Emb}(f,\prec_{e},\prec_{T})$ is false after all. Contradiction. ∎
It follows from the claims that $T\nvdash\mathsf{RFN}_{\Sigma^{1}_{1}}(T)$, which completes the proof of Theorem 2.7.